Shift Summary for Eve shift, 06/30/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Wed, 30 Jun 1999 22:21:46 -0700 (MST)

Shift Summary for Eve shift, 06/30/99

On Shift: Albert Erwin, Sydney Taegar

Hours of beam delivered at all intensities: 8.3 hrs
Hours of beam delivered at nominal intensity: 8.3 hrs
Hours of E832 data (special runs not included): 7.5 hrs
Hours of special runs (laser scans, muon runs...): 0.5 hrs

Accounting of lost time:

Loading new tapes
KFRENDs

========================================================================
Runs ended this shift:

Run 13774 - 11.5M events - BEAM_832
Run 13775 - 10.6M events - BEAM_832

Tapes brought to FCC this shift:

QKC067-069
QKS067-069
QKK067-069
QKB023

========================================================================
Accesses during shift:

NONE

========================================================================
Overview of shift:

Extremely smooth E832 datataking at 5E12 for both the
experiment and the accelerator.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13774 15:30 Shift begins with BEAM_832 run in progress.

15:40 Sasha L connects new LN2 dewar to TRD radiators.
140 KTEVANA errors -> KFREND.

151 15:49 KTEVANA errors -> KFREND.

172 16:18 KTEVANA errors -> KFREND.

175 16:21 KTEVANA errors -> KFREND.

223 17:31 No beam.
CH13: "LINAC Klystron 1 trip - Investigating."

236 17:41 Beam returns.

18:00 Gas checklist completed.

18:30 Tapes fill up, run ends.
Unload tapes QK*67-69, QKB023.
Load tapes QK*70-72, QKB024.

13775 18:45 Start BEAM_832 run.

0 18:50 KTEVANA ERRORS -> KFREND.

18:54 MU3 DYC busy -> KFREND.

19:10 DC histogram screen missing -> fix_disp.

22:00 Sandrine shows up early -- does extensive
research on Grand Canyon hiking.

23:35 Leo shows up for owl shift. Begins by
demanding coffee.

23:40 Stop run to take MUON run.
Switch NM4AN polarity to POS.
Fix BEAM_MUON dest file to point to
correct tape drive.

13776 23:54 Start MUON run.

24:00 Stop MUON run.
Vault tapes QK*67-69, QKB023.
Hand off to owl shift.

========================================================================
Follow up Notes on Previous Shift

Kudos to Hogan for alleviating the "Coffee Critical" situation.

========================================================================
Items to Follow up on Next Shift

E832 datataking.
Take LASER_SCAN if down for ~1 hour.